# NexusFlow: Unifying Disparate Tasks under Partial Supervision via Invertible Flow Networks

### v1

@@ 1 Introduction
Learning multiple tasks simultaneously via Multi-Task Learning (MTL) is a powerful paradigm for improving model efficiency and generalization while avoiding redundant training [xu2018pad, zhang2019pattern, vandenhende2020mti, li2024region]. However, its real-world applicability is often limited by the prohibitive cost of acquiring exhaustive annotations for every task, particularly in vision-heavy domains. In practice, datasets frequently lack labels for some tasks, or contain incomplete and unreliable annotations, motivating the crucial research direction of Partially Supervised Multi-Task Learning (PS-MTL) [li2022learning].
Significant progress has been made when tasks are homogeneous dense predictions, thanks to their natural interdependence [zamir2018taskonomy, zamir2020robust]. This has enabled joint training in both partially and fully supervised settings for tasks such as semantic segmentation, depth estimation, and surface normals estimation [eigen2014depth, he2017mask, chen2018gradnorm, poggi2020uncertainty, zhang2020uc]. Yet, when tasks are structurally disparate (e.g., one requiring dense pixel-wise labels and another producing sparse instance-level outputs), the challenge becomes far greater and has received very limited attention.
The challenge becomes even more pronounced under realistic supervision patterns. Existing PS-MTL methods typically simulate missing labels by randomly masking task annotations [li2024region], resulting in a controlled but overly simplified setup [ye2024diffusionmtl]. In practice, however, annotations for different tasks are often collected in disjoint domains (e.g., distinct geographic regions or capture environments). This creates a strong decoupling between task type and data domain, introducing substantial domain shift on top of structural disparity. Such domain-partitioned, heterogeneous-task conditions are common in real-world multi-task systems but remain largely unexplored. To the best of our knowledge, no prior work systematically addresses PS-MTL in this challenging and practically important regime.
@@
Inspired by flow-based invertible models [Dinh2014NICE, DinhSB17:realnvp, ardizzone2018analyzing, behrmann2019invertible, kobyzev2020normalizing], NexusFlow inserts, for each task, a dimensionality reduction module followed by invertible affine coupling layers (Figure 1). These coupling layers are bijective: they preserve all task-relevant information while enabling flexible transformations into a shared canonical space. This invertibility is essential; unlike conventional CNN-based alignment modules, the coupling layers prevent representational collapse beyond the reduced MLP embedding and allow structurally different tasks (e.g., sparse instance sets vs. dense geometric fields) to be aligned without sacrificing expressive capacity. By explicitly minimizing distributional discrepancies within this canonical space, NexusFlow encourages the model to discover cross-task feature representations that generalize even when each task is supervised in disjoint domains. This leads to a unified latent space that is more robust to both structural disparity and domain shift.
We validate NexusFlow in two complementary settings: (1) a structurally disparate, domain-partitioned two-task perception system on the nuScenes dataset [caesar2020nuscenes], involving multi-object tracking and map reconstruction under PS-MTL; and (2) a homogeneous three-task PS-MTL setup on NYUv2 [silberman2012indoor], covering dense prediction tasks of semantic segmentation, depth estimation, and surface normal estimation. Across both scenarios, NexusFlow not only delivers substantial gains over strong baselines but also surpasses baselines integrated with prior state-of-the-art PS-MTL modules developed for homogeneous tasks with random label masking, demonstrating strong generalization and broad applicability.
@@ 2 Related Work
Partially Supervised Multi-Task Learning. Partially Supervised Multi-Task Learning (PS-MTL) tackles the challenge of learning multiple tasks when only a subset has labels per sample. Early approaches showed success mainly in homogeneous dense prediction tasks (e.g., semantic segmentation, depth estimation), using strategies like consistency regularization, pseudo-labeling, or adversarial training. For example, adversarial discriminators have been used to align distributions across partially labeled datasets [wang2022semi], while consistency-based methods either employ model augmentation to provide pseudo-supervision [spinola2023knowledge] or regularize cross-task relationships in a shared space [li2022learning]. Pseudo-labeling has also been advanced via hierarchical task tokens for “label discovery,” enabling dense supervision for unlabeled tasks [zhang2024multi].
@@
Despite these advances, such techniques remain tailored to homogeneous tasks and are not directly applicable to structurally heterogeneous settings like object tracking versus map segmentation. Even recent unified frameworks for autonomous driving [hu2023planning, huang2023fuller] assume full supervision, leaving unresolved the realistic case of partially annotated, structurally disparate tasks—where annotation costs are prohibitive [dulac2021challenges, vettoruzzo2024advances, dai2025language]. A central challenge is learning shared representations without negative transfer. Architectural methods explore optimal layer sharing [ruder2019latent], while alignment-based methods project features into a common latent space. Early pairwise approaches [li2022learning] struggled with quadratic complexity, motivating more scalable solutions like JTR [nishi2024joint], which stacks all predictions into a unified joint-task space, or StableMTL [cao2025stablemtl], which applies a unified latent loss with efficient 1-to-N attention.
@@
Beyond scalability, newer methods seek richer cross-task knowledge transfer. Region-aware strategies use SAM [kirillov2023segment] to detect local regions and model their features as Gaussian distributions, enabling fine-grained alignment [li2024region, lin2024loss]. DiffusionMTL [ye2024diffusionmtl] reframes partially labeled outputs as noisy predictions and refines them through a denoising diffusion process with multi-task conditioning. These works highlight the trend towards more sophisticated, scalable PS-MTL approaches, but none address the harder setting of structurally disparate tasks under domain-partitioned supervision, which is the focus of our work.
@@
Multi-Task Learning for Autonomous Driving Perception. Autonomous driving perception has shifted from modular pipelines (separate detection, tracking, and mapping) to integrated multi-task learning (MTL) frameworks [wang2025generative, wang2025uniocc, gao2025safecoop], motivated by efficiency, performance, and shared representations [xu2022fpcc, luo2025v2x, xing2024autotrust, xing2025openemma].
@@
A major breakthrough is end-to-end models that unify perception, prediction, and planning. UniAD [hu2023planning] pioneered this direction with a Bird’s-Eye-View (BEV) representation and Transformer-based modules for tracking (TrackFormer), mapping (MapFormer), motion prediction (MotionFormer), and occupancy prediction (OccFormer), all linked by a query-based mechanism. Building on this, GenAD [zheng2024genad] models the scene generatively for joint prediction and planning, while DriveTransformer [jia2025drivetransformer] parallelizes tasks via shared attention for scalability. VAD [jiang2023vad] instead uses a fully vectorized scene representation, improving efficiency and reducing collision rates.
@@
A key challenge is domain generalization, as perception models face shifts in weather, lighting, sensors, and geography.  [wang2021generalizing] highlights this as central to real-world ML. For BEV-based systems,  [wang2023towards] analyzes domain gaps in 3D detection and proposes robust depth learning.  [jiang2024bev] introduces DA-BEV for unsupervised adaptation, combining image-view and BEV features.  [chang2024unified] unifies domain generalization and adaptation with multi-view overlap depth constraints. After all, the cost of annotation remains prohibitive: datasets are often partially labeled. Li  [li2022learning] addresses this by mapping task pairs into a joint space to enable sharing under incomplete supervision, a critical issue for large-scale autonomous driving systems where exhaustive labels are impractical.
@@ 3.1 Preliminaries
  • Domain-Partitioned Supervision: This is a far more challenging scenario than random masking. Supervision is assigned according to geographically distinct subsets of data. Specifically, in our experiments on the nuScenes dataset [caesar2020nuscenes], one task is annotated only in Boston scenes, while the other is annotated only in Singapore scenes. This means for any given sample, the mask is strict, and is perfectly correlated with the input’s geographic domain. This creates a substantial domain gap between supervision sources, which, when combined with structural disparity, dramatically increases the difficulty of knowledge transfer. This setup also applies to indoor domains with domain gaps between regionally distinct subsets.
@@ 3.2 NexusFlow
(2) Invertible Transformation. Each embedding is transformed by an affine coupling layer $c_{i}(\cdot)$ from normalizing flows [Dinh2014NICE]: $z_{i}=c_{i}\!\left(h^{\prime}_{i}\right)$. Given an embedding $h^{\prime}_{i}$ split into $(h_{i}^{{}^{\prime}1},h_{i}^{{}^{\prime}2})$, the layer performs:
@@
where $s(\cdot)$ and $t(\cdot)$ are small MLPs. This bijective transformation maps features into a canonical latent space while preserving information content.
@@
$\mathcal{L}_{\mathrm{align(pair)}}=\sum_{i<j}\|z^{i}-z^{j}\|_{2}.$ (3)
@@
$\mathcal{L}_{\mathrm{align(center)}}=\sum_{i=1}^{n}\|z^{i}-\bar{z}\|_{2}.$ (4)
Since the pairwise formulation requires $O(n^{2})$ pairwise comparisons, whereas the center-based formulation uses only $O(n)$ terms, we adopt the center-based objective for its substantially higher efficiency and simpler gradient structure. For completeness, we also compare both variants in Table 4.
@@ 3.3 Theoretical Analysis
Let $h^{\prime}_{1},h^{\prime}_{2}\in\mathbb{R}^{N}$ be the compact features of two tasks $t_{1}$ and $t_{2}$ passed into their coupling layers $c_{1}$ and $c_{2}$. Assume their inverse transformations $c_{1}^{-1}$ and $c_{2}^{-1}$ are $L$-Lipschitz continuous with constant $L$ [virmaux2018lipschitz, gouk2021regularisation]. Then the L2 distance between the original features is upper-bounded by
@@ 3.4 Practical Analysis
Building upon our theoretical guarantee, we provide a practical analysis to empirically verify the core mechanism of NexusFlow. We hypothesize that by explicitly aligning the latent feature distributions, our method forges a unified and more effective representation space for knowledge transfer. To validate this, we investigate two key properties of the learned representations: their alignment and intrinsic dimensionality. We assess alignment both qualitatively through t-SNE visualizations [maaten2008visualizing, cai2022theoretical] and quantitatively using the Maximum Mean Discrepancy (MMD) metric [gretton2012kernel, lin2023hyperbolic]. We then analyze their intrinsic dimensionality via Principal Component Analysis (PCA) [abdi2010principal, del2021effective] to ensure the unified representation is also complex enough to serve both disparate tasks.
@@
We first evaluate the alignment of the feature distributions. For a qualitative understanding, we employ t-SNE [maaten2008visualizing, cai2022theoretical] to visualize two (or three) sets of high-dimensional features: the latent variables produced by our NexusFlow module. As illustrated in Figure 2 and Figure 3, the visualization is revealing. The critical distinction shows the latent variables after the our designed surrogate networks. Here, the features from the baseline methods still exhibit a clear distributional shift. In contrast, the latent variables from our NexusFlow model become substantially intermingled, providing visual evidence that our method creates a more unified manifold for knowledge transfer.
This qualitative observation is substantiated by our quantitative analysis using the Maximum Mean Discrepancy (MMD). MMD is a standard metric for measuring the distance between two distributions, where a lower score signifies greater similarity [gretton2012kernel, yue2024understanding, chen2024fair]. We compute the MMD between the intermediate features(after NexusFlow module) from mapping-annotated samples and tracking-annotated samples on nuScenes dataset, we also compute the MMD among segmentation-annotated samples, depth-annotated samples and normal-annotated samples on NYU-V2 datasets. As shown in Table 1, on both datasets, NexusFlow module leads to a significant reduction in the MMD score compared to all baselines. This quantitatively validates that NexusFlow effectively reduces the distributional distance between the disparate tasks, creating the statistical foundation for knowledge transfer.
@@
Beyond showing the distributions are closer, we analyze the feature distributions’ intrinsic complexity using Principal Component Analysis (PCA). Our analysis is guided by the premise that a representation with a slower decay in its eigenvalue magnitudes holds more informative dimensions, making it more suitable for complex, multi-tasks [goel2017eigenvalue, ansuini2019intrinsic, kim2023vne, zhang2025gps]. As shown in the Scree plot [zhu2006automatic] in Figure 4, the features from the baseline and a variant of our model where we ablate the invertible layer (‘Ours (w/o inv)‘) exhibit rapid eigenvalue decay, suggesting their representations are more compressible and may lose nuanced information. In contrast, the eigenvalue of our full NexusFlow model decays at a slower rate. This provides evidence that our alignment process forges a more complex and information-rich feature space, one that retains the high dimensionality necessary to effectively serve two structurally disparate tasks.
@@ 4.1 Core Challenge on nuScenes
Dataset and PS-MTL Protocol. We use nuScenes [caesar2020nuscenes], a large-scale autonomous driving dataset captured in Boston and Singapore. It provides annotations for our two disparate tasks: multi-object tracking ($\mathcal{T}_{track}$) and map reconstruction ($\mathcal{T}_{map}$). To rigorously evaluate the performance under an explicit domain gap, we design a specific data protocol: ground-truth annotations for the mapping task are only provided for scenes in Boston, while annotations for the tracking task are only provided for scenes in Singapore. This challenging, geographically-partitioned supervision protocol is applied consistently across all evaluated methods.
Baselines and Architecture. To ensure a fair comparison, all methods are built upon the SOTA UniAD [hu2023planning] architecture as the backbone. We compare against several methods: (1) Oracle, trained with full supervision for both tasks from both cities. (2) Baseline, representing the standard UniAD model trained on our domain-partitioned data. (3) Two re-implemented PS-MTL methods, MTPSL [li2022learning] and JTR [nishi2024joint], whose core mechanisms we adapt and integrate into the UniAD framework.
@@
Oracle (fully supervised) 0.323 1.328 0.431 696 31.4 71.42 34.4 21.3
Baseline [hu2023planning] 0.289 1.488 0.362 1025 27.1 62.7 22.6 14.1
MTPSL [li2022learning] 0.255 1.504 0.321 1089 27.0 59.6 21.7 11.5
JTR [nishi2024joint] 0.197 1.547 0.317 774 25.1 57.6 21.0 12.1
@@
Implementation Details. For the nuScenes experiments, all methods are trained from scratch following the default setting in UniAD [hu2023planning] on four NVIDIA A100 80G GPUs. Our proposed NexusFlow module, implemented with $N=4$ invertible coupling layers, is inserted between the shared BEV encoder and the two disparate task decoders to align their latent feature distributions.
Figure 5: Visualization results. We show results for multi-object tracking and online mapping tasks in surround-view images and BEV from left to right. We highlight the difference with red dashed circle.
Quantitative Results. We present our results in Table 2, with the main metric for each task highlighted in gray. While training is strictly conducted under our scenario-based partial supervision protocol, evaluation is performed on the complete validation set, including scenes from both Boston and Singapore. Our proposed NexusFlow significantly outperforms all partially supervised baselines and, remarkably, achieves performance competitive with the fully supervised Oracle.
  • Multi-Object Tracking: For the tracking task (left side of the table), NexusFlow sets a new SOTA for this challenging PS-MTL setting. It surpasses MTPSL [li2022learning] by a large margin of +7.4% AMOTA and improves upon the standard Baseline [hu2023planning] by +4.0% AMOTA. Furthermore, our method achieves the lowest ID Switch score, demonstrating a superior ability to maintain temporal consistency for each tracklet.
  • Online Mapping: For the mapping task (right side of the table), our method shows substantial gains in segmenting crucial road elements. Notably, it outperforms the Baseline [hu2023planning] and MTPSL [li2022learning] on lanes by over +10% IoU, a critical component for safe downstream motion planning. This result confirms that effective knowledge was transferred from the tracking-annotated domain (Singapore) to enhance the mapping performance.
Meanwhile, we observe that prior state-of-the-art PS-MTL methods, MTPSL [li2022learning] and JTR [nishi2024joint], designed for homogeneous dense prediction tasks can even degrade baseline performance in our setting. This highlights that the more realistic scenario of structurally disparate, domain-partitioned partial labels can break methodologies tailored only for small-scale, homogeneous multi-task learning.
Qualitative Results. These quantitative results are corroborated by our qualitative analysis in Figure 5. We observe that our model is able to detect objects more accurately. A full video comparison is provided in the supplementary material.

### v2

@@ 1 Introduction
Learning multiple tasks simultaneously via Multi-Task Learning (MTL) is a powerful paradigm for improving model efficiency and generalization while avoiding redundant training compared with single task [53, 61, 42, 54, 32, 56, 30, 50]. However, its real-world applicability is often limited by the prohibitive cost of acquiring exhaustive annotations for every task, particularly in vision-heavy domains. In practice, datasets frequently lack labels for some tasks, or contain incomplete and unreliable annotations, motivating the crucial research direction of Partially Supervised Multi-Task Learning (PS-MTL) [31].
Significant progress has been made when tasks are homogeneous dense predictions, thanks to their natural interdependence [57, 58]. This has enabled joint training in both partially and fully supervised settings for tasks such as semantic segmentation, depth estimation, and surface normals estimation [16, 21, 10, 38, 59]. Yet, when tasks are structurally disparate (e.g., one requiring dense pixel-wise labels and another producing sparse instance-level outputs), the challenge becomes far greater and has received very limited attention.
The challenge becomes even more pronounced under realistic supervision patterns. Existing PS-MTL methods typically simulate missing labels by randomly masking task annotations [30], resulting in a controlled but overly simplified setup [55]. In practice, however, annotations for different tasks are often collected in disjoint domains (e.g., distinct geographic regions or capture environments). This creates a strong decoupling between task type and data domain, introducing substantial domain shift on top of structural disparity. Such domain-partitioned, heterogeneous-task conditions are common in real-world multi-task systems but remain largely unexplored. To the best of our knowledge, no prior work systematically addresses PS-MTL in this challenging and practically important regime.
@@
Inspired by flow-based invertible models [13, 14, 3, 4, 29], NexusFlow inserts, for each task, a dimensionality reduction module followed by invertible affine coupling layers (Figure 1). These coupling layers are bijective: they preserve all task-relevant information while enabling flexible transformations into a shared canonical space. This invertibility is essential; unlike conventional CNN-based alignment modules, the coupling layers prevent representational collapse beyond the reduced MLP embedding and allow structurally different tasks (e.g., sparse instance sets vs. dense geometric fields) to be aligned without sacrificing expressive capacity. By explicitly minimizing distributional discrepancies within this canonical space, NexusFlow encourages the model to discover cross-task feature representations that generalize even when each task is supervised in disjoint domains. This leads to a unified latent space that is more robust to both structural disparity and domain shift.
We validate NexusFlow in two complementary settings: (1) a structurally disparate, domain-partitioned two-task perception system on the nuScenes dataset [5], involving multi-object tracking and map reconstruction under PS-MTL; and (2) a homogeneous three-task PS-MTL setup on NYUv2 [40], covering dense prediction tasks of semantic segmentation, depth estimation, and surface normal estimation. Across both scenarios, NexusFlow not only delivers substantial gains over strong baselines but also surpasses baselines integrated with prior state-of-the-art PS-MTL modules developed for homogeneous tasks with random label masking, demonstrating strong generalization and broad applicability.
@@ 2 Related Work
Partially Supervised Multi-Task Learning. Partially Supervised Multi-Task Learning (PS-MTL) tackles the challenge of learning multiple tasks when only a subset has labels per sample. Early approaches showed success mainly in homogeneous dense prediction tasks (e.g., semantic segmentation, depth estimation), using strategies like consistency regularization, pseudo-labeling, or adversarial training. For example, adversarial discriminators have been used to align distributions across partially labeled datasets [47], while consistency-based methods either employ model augmentation to provide pseudo-supervision [41] or regularize cross-task relationships in a shared space [31]. Pseudo-labeling has also been advanced via hierarchical task tokens for “label discovery,” enabling dense supervision for unlabeled tasks [60].
@@
Despite these advances, such techniques remain tailored to homogeneous tasks and are not directly applicable to structurally heterogeneous settings like object tracking versus map segmentation. Even recent unified frameworks for autonomous driving [22, 23, 48, 52] assume full supervision, leaving unresolved the realistic case of partially annotated, structurally disparate tasks—where annotation costs are prohibitive [15, 33, 43, 62, 11]. A central challenge is learning shared representations without negative transfer. Architectural methods explore optimal layer sharing [39], while alignment-based methods project features into a common latent space. Early pairwise approaches [31] struggled with quadratic complexity, motivating more scalable solutions like JTR [37], which stacks all predictions into a unified joint-task space, or StableMTL [7], which applies a unified latent loss with efficient 1-to-N attention.
@@
Beyond scalability, newer methods seek richer cross-task knowledge transfer. Region-aware strategies use SAM [28] to detect local regions and model their features as Gaussian distributions, enabling fine-grained alignment [30]. DiffusionMTL [55] reframes partially labeled outputs as noisy predictions and refines them through a denoising diffusion process with multi-task conditioning. These works highlight the trend towards more sophisticated, scalable PS-MTL approaches, but none address the harder setting of structurally disparate tasks under domain-partitioned supervision, which is the focus of our work.
@@
Multi-Task Learning for Autonomous Driving Perception. Autonomous driving perception has shifted from modular pipelines (separate detection, tracking, and mapping) to integrated multi-task learning (MTL) frameworks [49, 48, 17], motivated by efficiency, performance, and shared representations [54, 35, 51, 52].
@@
A major breakthrough is end-to-end models that unify perception, prediction, and planning. UniAD [22] pioneered this direction with a Bird’s-Eye-View (BEV) representation and Transformer-based modules for tracking (TrackFormer), mapping (MapFormer), motion prediction (MotionFormer), and occupancy prediction (OccFormer), all linked by a query-based mechanism. Building on this, GenAD [63] models the scene generatively for joint prediction and planning, while DriveTransformer [24] parallelizes tasks via shared attention for scalability. VAD [25] instead uses a fully vectorized scene representation, improving efficiency and reducing collision rates.
@@
A key challenge is domain generalization, as perception models face shifts in weather, lighting, sensors, and geography.  [45] highlights this as central to real-world ML. For BEV-based systems,  [46] analyzes domain gaps in 3D detection and proposes robust depth learning.  [26] introduces DA-BEV for unsupervised adaptation, combining image-view and BEV features.  [8] unifies domain generalization and adaptation with multi-view overlap depth constraints. After all, the cost of annotation remains prohibitive: datasets are often partially labeled. Li  [31] addresses this by mapping task pairs into a joint space to enable sharing under incomplete supervision, a critical issue for large-scale autonomous driving systems where exhaustive labels are impractical.
@@ 3.1 Preliminaries
  • Domain-Partitioned Supervision: This is a far more challenging scenario than random masking. Supervision is assigned according to geographically distinct subsets of data. Specifically, in our experiments on the nuScenes dataset [5], one task is annotated only in Boston scenes, while the other is annotated only in Singapore scenes. This means for any given sample, the mask is strict, and is perfectly correlated with the input’s geographic domain. This creates a substantial domain gap between supervision sources, which, when combined with structural disparity, dramatically increases the difficulty of knowledge transfer. This setup also applies to indoor domains with domain gaps between regionally distinct subsets.
@@ 3.2 NexusFlow
(2) Invertible Transformation. Each embedding is transformed by an affine coupling layer $c_{i}(\cdot)$ from normalizing flows [13]: $z_{i}=c_{i}\!\left(h^{\prime}_{i}\right)$. We use the standard affine coupling transform introduced in RealNVP [14]. Given an embedding $h^{\prime}_{i}$ split into $(h_{i}^{{}^{\prime}1},h_{i}^{{}^{\prime}2})$, the layer performs:
@@
where $s(\cdot)$ and $t(\cdot)$ are small MLPs. Due to the design of the coupling transform, its inverse admits a closed-form expression that directly reuses the forward $s(\cdot)$ or $t(\cdot)$ functions, without requiring inversion of these MLPs [14]. This bijective transformation maps features into a canonical latent space while preserving information content.
@@
$\mathcal{L}_{\mathrm{align(pair)}}=\sum_{i<j}\|z^{i}-z^{j}\|_{2}^{2}.$ (3)
@@
$\mathcal{L}_{\mathrm{align(center)}}=\sum_{i=1}^{n}\|z^{i}-\bar{z}\|_{2}^{2}.$ (4)
The pairwise formulation requires $O(n^{2})$ pairwise comparisons, whereas the center-based formulation uses only $O(n)$ terms, we adopt the center-based objective for its substantially higher efficiency and simpler gradient structure. For completeness, we also compare both variants in Table 4.
@@ 3.3 Theoretical Analysis
Let $h^{\prime}_{1},h^{\prime}_{2}\in\mathbb{R}^{N}$ be the compact features of two tasks $t_{1}$ and $t_{2}$ passed into their coupling layers $c_{1}$ and $c_{2}$. Assume their inverse transformations $c_{1}^{-1}$ and $c_{2}^{-1}$ are $L$-Lipschitz continuous with constant $L$ [44, 19]. Then the L2 distance between the original features is upper-bounded by
@@
More specifically, when two tasks are aligned in the latent space via invertible coupling layers, this alignment provably transfers back to the original task-specific feature spaces, up to a controllable error.
@@ 3.4 Practical Analysis
Building upon our theoretical guarantee, we provide a practical analysis to empirically verify the core mechanism of NexusFlow. We hypothesize that by explicitly aligning the latent feature distributions, our method forges a unified and more effective representation space for knowledge transfer. To validate this, we investigate two key properties of the learned representations: their alignment and intrinsic dimensionality. We assess alignment both qualitatively through t-SNE visualizations [36, 6] and quantitatively using the Maximum Mean Discrepancy (MMD) metric [20, 62]. We then analyze their intrinsic dimensionality via Principal Component Analysis (PCA) [1, 12] to ensure the unified representation is also complex enough to serve both disparate tasks.
@@
We first evaluate the alignment of the feature distributions. For a qualitative understanding, we employ t-SNE [36, 6] to visualize two (or three) sets of high-dimensional features: the latent variables produced by our NexusFlow module. As illustrated in Figure 2 and Figure 3, the visualization is revealing. The critical distinction shows the latent variables after the our designed surrogate networks. Here, the features from the baseline methods still exhibit a clear distributional shift. In contrast, the latent variables from our NexusFlow model become substantially intermingled, providing visual evidence that our method creates a more unified manifold for knowledge transfer.
This qualitative observation is substantiated by our quantitative analysis using the Maximum Mean Discrepancy (MMD). MMD is a standard metric for measuring the distance between two distributions, where a lower score signifies greater similarity [20, 56, 9]. We compute the MMD between the intermediate features(after NexusFlow module) from mapping-annotated samples and tracking-annotated samples on nuScenes dataset, we also compute the MMD among segmentation-annotated samples, depth-annotated samples and normal-annotated samples on NYU-V2 datasets. As shown in Table 1, on both datasets, NexusFlow module leads to a significant reduction in the MMD score compared to all baselines. This quantitatively validates that NexusFlow effectively reduces the distributional distance between the disparate tasks, creating the statistical foundation for knowledge transfer.
@@
Beyond showing the distributions are closer, we analyze the feature distributions’ intrinsic complexity using Principal Component Analysis (PCA). Our analysis is guided by the premise that a representation with a slower decay in its eigenvalue magnitudes holds more informative dimensions, making it more suitable for complex, multi-tasks [18, 2, 27, 62]. As shown in the Scree plot [64] in Figure 4, the features from the baseline and a variant of our model where we ablate the invertible layer (‘Ours (w/o inv)‘) exhibit rapid eigenvalue decay, suggesting their representations are more compressible and may lose nuanced information. In contrast, the eigenvalue of our full NexusFlow model decays at a slower rate. This provides evidence that our alignment process forges a more complex and information-rich feature space, one that retains the high dimensionality necessary to effectively serve two structurally disparate tasks.
@@ 4.1 Core Challenge on nuScenes
Dataset and PS-MTL Protocol. We use nuScenes [5], a large-scale autonomous driving dataset captured in Boston and Singapore. It provides annotations for our two disparate tasks: multi-object tracking ($\mathcal{T}_{track}$) and map reconstruction ($\mathcal{T}_{map}$). To rigorously evaluate the performance under an explicit domain gap, we design a specific data protocol: ground-truth annotations for the mapping task are only provided for scenes in Boston, while annotations for the tracking task are only provided for scenes in Singapore. This challenging, geographically-partitioned supervision protocol is applied consistently across all evaluated methods.
Baselines and Architecture. To ensure a fair comparison, all methods are built upon the SOTA UniAD [22] architecture as the backbone. We compare against several methods: (1) Full-supervision , trained with full supervision for both tasks from both cities. (2) Baseline, representing the standard UniAD model trained on our domain-partitioned data. (3) Two re-implemented PS-MTL methods, MTPSL [31] and JTR [37], whose core mechanisms we adapt and integrate into the UniAD framework.
@@
Full-supervision 0.323 1.328 0.431 696 31.4 71.42 34.4 21.3
Baseline [22] 0.289 1.488 0.362 1025 27.1 62.7 22.6 14.1
MTPSL [31] 0.255 1.504 0.321 1089 27.0 59.6 21.7 11.5
JTR [37] 0.197 1.547 0.317 774 25.1 57.6 21.0 12.1
@@
Implementation Details. For the nuScenes experiments, all methods are trained from scratch following the default setting in UniAD [22] on four NVIDIA A100 80G GPUs. Our proposed NexusFlow module, implemented with $N=4$ invertible coupling layers, is inserted between the shared BEV encoder and the two disparate task decoders to align their latent feature distributions.
Figure 5: We show visualization results for multi-object tracking and online mapping tasks in surround-view images and BEV from left to right. We highlight the difference with red dashed circle.
Quantitative Results. We present our results in Table 2, the main metric for each task highlighted in gray. Training is strictly conducted under our scenario-based partial supervision protocol, evaluation is performed on the complete validation set, including scenes from both Boston and Singapore. Our proposed NexusFlow significantly outperforms all partially supervised baselines and, remarkably, achieves performance competitive with the Full-supervision.
  • Multi-Object Tracking: For the tracking task (left side of the table), NexusFlow sets a new SOTA for this challenging PS-MTL setting. It surpasses MTPSL [31] by a large margin of +7.4% AMOTA and improves upon the standard Baseline [22] by +4.0% AMOTA. Furthermore, our method achieves the lowest ID Switch score, demonstrating a superior ability to maintain temporal consistency for each tracklet.
  • Online Mapping: For the mapping task (right side of the table), our method shows substantial gains in segmenting crucial road elements. Notably, it outperforms the Baseline [22] and MTPSL [31] on lanes by over +10% IoU, a critical component for safe downstream motion planning. This result confirms that effective knowledge was transferred from the tracking-annotated domain (Singapore) to enhance the mapping performance.
Meanwhile, we observe that prior state-of-the-art PS-MTL methods, MTPSL [31] and JTR [37], designed for homogeneous dense prediction tasks can even degrade baseline performance in our setting. This highlights that the more realistic scenario of structurally disparate, domain-partitioned partial labels can break methodologies tailored only for small-scale, homogeneous multi-task learning.
Qualitative Results. These results are corroborated by our qualitative analysis in Figure 5. We observe that our model is able to detect objects more accurately. A full video comparison is provided in the supplementary material.